OpenStudy (anonymous):

What is 2 3/4 - 1/2? A. 1 1/4 B. 2 1/4 C. 3 D. -2 1/4

OpenStudy (anonymous):

u want to c the work?

OpenStudy (anonymous):

answer is B.

OpenStudy (anonymous):

yes thanks!!! :D

OpenStudy (anonymous):

make 2 3/4 into a n improper fraction: 11/4 now make 1/2 nd 11/4 have the same denominator: gcd=4 so 1/2 * 2/2 = 2/4 nopw subtract 11/4-2/4 11-2=9 so: 9/4 make 9/4 into a mixed number: 2 1/4

OpenStudy (anonymous):

do u get it?